20 The risk-free rate is 3.96% and the market risk premium is 9.00%. A stock with...

20

The risk-free rate is 3.96% and the market risk premium is 9.00%. A stock with a β of 1.19 just paid a dividend of $2.84. The dividend is expected to grow at 20.37% for three years and then grow at 4.84% forever. What is the value of the stock?

Answer format: Currency: Round to: 2 decimal places.

#3 The risk-free rate is 3.70% and the market risk premium is 7.42%. A stock with a β of 1.32 just paid a dividend of $1.09. The dividend is expected to grow at 21.86% for five years and then grow at 4.23% forever. What is the value of the stock?

Answer format: Currency: Round to: 2 decimal places.

#4 Caspian Sea Drinks needs to raise $24.00 million by issuing additional shares of stock. If the market estimates CSD will pay a dividend of $2.42 next year, which will grow at 4.46% forever and the cost of equity to be 10.81%, then how many shares of stock must CSD sell?

Answer format: Number: Round to: 0 decimal places.

#5 Suppose the risk-free rate is 2.27% and an analyst assumes a market risk premium of 7.89%. Firm A just paid a dividend of $1.14 per share. The analyst estimates the β of Firm A to be 1.45 and estimates the dividend growth rate to be 4.44% forever. Firm A has 250.00 million shares outstanding. Firm B just paid a dividend of $1.77 per share. The analyst estimates the β of Firm B to be 0.82 and believes that dividends will grow at 2.04% forever. Firm B has 180.00 million shares outstanding. What is the value of Firm A?

Answer format: Currency: Round to: 2 decimal places.

#6 Suppose the risk-free rate is 1.25% and an analyst assumes a market risk premium of 6.18%. Firm A just paid a dividend of $1.18 per share. The analyst estimates the β of Firm A to be 1.50 and estimates the dividend growth rate to be 4.99% forever. Firm A has 284.00 million shares outstanding. Firm B just paid a dividend of $1.72 per share. The analyst estimates the β of Firm B to be 0.89 and believes that dividends will grow at 2.85% forever. Firm B has 192.00 million shares outstanding. What is the value of Firm B?
Answer #1

1.
=2.84*((1+20.37%)/(1+3.96%+1.19*9%))^1+2.84*((1+20.37%)/(1+3.96%+1.19*9%))^2+2.84*((1+20.37%)/(1+3.96%+1.19*9%))^3+2.84*((1+20.37%)/(1+3.96%+1.19*9%))^3*(1+4.84%)/(3.96%+1.19*9%-4.84%)=44.4300410203936

2.
=1.09*((1+21.86%)/(1+3.7%+1.32*7.42%))^1+1.09*((1+21.86%)/(1+3.7%+1.32*7.42%))^2+1.09*((1+21.86%)/(1+3.7%+1.32*7.42%))^3+1.09*((1+21.86%)/(1+3.7%+1.32*7.42%))^4+1.09*((1+21.86%)/(1+3.7%+1.32*7.42%))^5+1.09*((1+21.86%)/(1+3.7%+1.32*7.42%))^5*1.0423/(3.7%+1.32*7.42%-4.23%)=24.2802350637436

3.
=24*10^6/(2.42/(10.81%-4.46%))=629752.066115703

4.
=1.14*1.044/(2.27%+1.45*7.89%-4.44%)=12.8381424950111